OpenStudy (anonymous):

The ages of trees in a forest are normally distributed with a mean of 25 years and a standard deviation of 4. Approximately what percent of the trees are between 20 and 30 years old? @Mathbreaker

OpenStudy (anonymous):

Calculate the z-scores for values = 20 and 30 z = (value - mean)/(standard deviation) look up cumulative normal distribution and find the percentage between these two z-scores.

OpenStudy (anonymous):

do i do 20-25 or 20-30-25? @douglaswinslowcooper

OpenStudy (anonymous):

find cumulative dnormal distribution for z=(20-mean)/ s.d. and for z = (30-man)/s.d. and the difference between these two cumulative percentages is the answer you want, the percentage between 20 and 30. If you had z=1 so P(1) = 84% and z=0, so P(0) = 50%, then you would have 84-50=34% between thse two z scores.